sketch the graph of the parabola
Find (f-g)(-4) when f(x)= x+2 and g(x) = x-3

Answer by stanbon(75887)

sketch the graph of the parabola
Find (f-g)(-4) when f(x)= x+2 and g(x) = x-3
------------
(f-g)(-4) means f(-4)-g(-4) = -2 - -7 = 5
-----------
There is no parabola to graph.
Cheers,
Stan H.